OpenStudy (anonymous):

Solve the equation 2 / 3 (x − 4) = 2x for x.

OpenStudy (aem9296):

\[\frac{ 2 }{ 3 } (x-4)=2x\] \[\frac{ 2 }{ 3}x - \frac{ 8 }{ 3} = 2x\] \[2x - 8 = 6x\] \[-8 = 4x\] \[x=-2\]

OpenStudy (anonymous):

thanks! <3